Please show work2.) Which is better for student learning - online HW or handwritten HW? Suppose students who do HW online are equally success
Answer #1

a) The null hypothesis is   Ho1-20    and the alternative hypothesis is   Ha 12 O .

(where   \mu_1    population average using online HW and   2    population average using handwritten HW.)

b) Type-I error:

In this problem type-I error is the error generated due to rejecting the null hypothesis based on the sample information i.e. concluding that the average success using online HW is different when actually the average success using online HW is not significantly different for the population.

Type-II error:

In this problem type-II error is the error generated due to accepting the null hypothesis based on the sample information i.e. concluding that the average success using online HW is not different when actually the average success using online HW is significantly different for the population.

c) (Assuming the two population standard deviations are the same because the two given sample standard deviations are same and we know that, sample standard deviations are the good unbiased estimators of population standard deviations)

The sample statistic for this test is   X T = 2 1 1 s S n1 n2 E ,

where   (n,-1)s (n2 -1) s2 308.038835 n1n2 2 X, Х   are the sample means   and   n_1,n_2    are the sample sizes of the respective samples.

Let the population standard deviation for both the samples is   \sigma .

(Assuming the samples are from   2 W(11,0    and   N(\mu_2,\sigma^2)    respectively.)

We know   X~ N(1,    and   X2~ N(42 n2

\therefore \bar{X_1}-\bar{X_2} \sim N\left ( \mu_1-\mu_2, \frac{\sigma^2}{n_1}+\frac{\sigma^2}{n_2} \right )

i.e.   \frac{(\bar{X_1}-\bar{X_2})-(\mu_1-\mu_2)}{\sigma\sqrt{\frac{1}{n_1}+\frac{1}{n_2}}} \sim N(0,1)

Now under   H_0 ,

\frac{\bar{X_1}-\bar{X_2}}{\sigma\sqrt{\frac{1}{n_1}+\frac{1}{n_2}}} \sim N(0,1)

And we know   \frac{(n_1-1)s_1^2}{\sigma^2} \sim \chi^2_{n_1-1}    and   \frac{(n_2-1)s_2^2}{\sigma^2} \sim \chi^2_{n_2-1}

So,   \frac{(n_1-1)s_1^2+(n_2-1)s_2^2}{\sigma^2} \sim \chi^2_{n_1+n_2-2} \Rightarrow \frac{(n_1+n_2-2)s^2}{\sigma^2}\sim \chi^2_{n_1+n_2-2}

So under   H_0 ,   \frac{\frac{\bar{X_1}-\bar{X_2}}{\sigma\sqrt{\frac{1}{n_1}+\frac{1}{n_2}}}}{ \sqrt{\frac{(n_1+n_2-2)s^2}{\sigma^2}/(n_1+n_2-2)}} \sim t_{n_1+n_2-2} \Rightarrow T \sim t_{n_1+n_2-2} .

So the sample statistic follows t distribution with degrees of freedom (n_1+n_2-2) .

Here the value of the sample statistic is   T=\frac{77.5-75.8}{\sqrt{308.038835}\times\sqrt{\frac{1}{60}+\frac{1}{45}}}=0.491172(approx.)

d) Here p-value is the probability of the sample statistic being more extreme than the calculated value of it, i.e. the probability of   T    being more extreme than the calculated value 0.491172.

The p-value of the test is,

2Min \left \{ P(T\leq0.491172),P(T\geq0.491172) \right \} =2P(T\geq0.491172)

=2\times 0.3121749=0.6243498(approx.)

T049 T O491172 CS Scanned with CamScanner

The above probability represents the shaded area is the picture.

e) In this test, we failed to reject the null hypothesis based on the sample information since the p-value of the test is greater than the given significance level   \alpha=0.05 .